# SMC Is All You Need: Parallel Strong Scaling

### v1

@@ 4.1 High-dimensional Gaussian cases
Consider a toy Gaussian inference example with $m$ observations ($y$) and $d$-dimensional parameter ($\theta$). Assume we have data $\mathcal{D}=\{y,X\}$, where $y\in\mathbb{R}^{m}$ and $X\in\mathbb{R}^{m\times d}$, and parameter $\theta\in\mathbb{R}^{d}$ connected by the following inverse problem:
@@
In the following tests, we compare pSMC-pCN with pCN performance in estimation of the posterior mean, $\mu=\mathbb{E}[\theta|\mathcal{D}]$. Let $X\in\mathbb{R}^{m\times d}$ be a randomly selected full rank matrix and $\sigma=0.01$. The observations are generated as
@@
The expected $-1$ convergence rate of pSMC-pCN and pCN are verified numerically in the left plots of figures in Appendix D.1. Note for the partially observed case, we have verified that convergence does happen if we run the chain for much much longer (not shown). Recall we defined the crossover point for fixed $M,N,J$ as the minimum number of nodes when pSMC has a lower MSE than MCMC. See Figure 1. We report the current MSE, the number of nodes for pSMC-pCN, and the time at the crossover points with pCN for the various different regimes in Table 2. The full convergence diagrams are given in the Appendix D.1. The improvement of pSMC is clearly the greatest in the wide case of partial observations $m<d$ for the Gaussian, suggesting that the method may be quite well-suited to small(er) data cases.
4.2 Quantum State Tomography problems
Quantum State Tomography (QST) is one application area where an accurate approximation of the Bayesian mean estimator is particularly valuable, but due to exponential scaling of the state-space as the number of qubits grows it quickly becomes computationally prohibitive [Lukens et al., 2020, 2021]. In a $Q$-qubit quantum system with $2-$level quantum information carriers, we intend to infer the $2^{Q}\times 2^{Q}$ density matrix $\rho(\theta)$ through the parameter $\theta\in\mathbb{R}^{d}$, where the $\rho$ is parameterized [^3] Details for the parameterization of $\rho(\theta)$ is given in Appendix B by $\theta\in\mathbb{R}^{d}$ with $d=2^{2(Q+1)}$ such that any value within $\theta$’s support returns a physical matrix. We consider using the Bayesian state estimation on $3^{Q}$ measurements with $2^{Q}$ outcomes for $S$ quantum states, respectively. There are then $m=6^{Q}$ total measurements for each quantum state, which can be described by a sequence of positive operators $\Lambda_{i}$ for $i\in\{1,...,m\}$ and corresponding observed counts $\mathcal{D}=\{c_{1},...,c_{m}\}$.
Following from Born’s rule, the likelihood is given by
Hence, the posterior distribution is,
$\pi(\theta)\propto\mathcal{L}(\theta)\pi_{0}(\theta),$ (15)
where $\pi_{0}$ is a standard Gaussian prior $N(0,I)$.
We compare pSMC-pCN (with $N=32$) and pCN for $Q=1,2,3$ qubits, i.e. $(d,m)=(16,6),(64,36),(256,216)$. The quantity of interest is $\varphi=\rho$. See Figures 7, 8 and 9. The expected $1/NR$ convergence rate for pSMC is verified in the left plots in the above figures. The right plots in those figures represent the strong scaling property in pSMC and the crossover point is shown in Table 3.
4.3 Bayesian Neural Network Examples
HMC has gained widespread acceptance and favour in high-dimensional models, like the neural networks. In this section, we compare pSMC-HMC with HMC using BNN examples with similar settings cited in Cobb and Jalaian [2021].
Suppose we have data set $\mathcal{D}=\{(x_{1},y_{1}),...,(x_{m},y_{m})\}$, where $x_{i}\in\mathbb{R}^{n}$ are inputs associated to output labels $y_{i}\in\mathsf{Y}$. Let $D\in\mathbb{N}$, $(n_{0},...,n_{D})\in\mathbb{N}^{D+1}$ with $n_{0}=n$ for the input layer. Let weights be $A_{i}\in\mathbb{R}^{n_{i}\times n_{i-1}}$ and biases be $b_{i}\in\mathbb{R}^{n_{i}}$ for $i\in\{1,...,D\}$, we denote $\theta:=((A_{1},b_{1}),...,(A_{D},b_{D}))$. The layer is defined by
@@ 4.3 Bayesian Neural Network Examples
In the regression example, we have $\mathsf{Y}=\mathbb{R}^{n_{D}}$ and can assume the data as follows, in which the data is defined in a continuous space,
$y_{i}=g(x_{i},\theta)+\nu_{i},\ \epsilon\sim N(0,\sigma^{2}I),$ (16)
where $i=\{1,...,m\}$. Then the likelihood for the regression problem (16) can be computed as
$\mathcal{L}(\theta)=\prod_{i=1}^{m}\exp(-\frac{1}{2\sigma^{2}}\left\lVert y_{i}-g(x_{i},\theta)\right\rVert^{2}).$
In the discrete data set, like the classification problems, we have $\mathsf{Y}=\{1,...,K\}$ and $n_{D}=K$, then we instead define the so-called softmax function as
@@
Now, the likelihood for the classification problem (17) can be computed as
where $\mathds{1}_{[y_{i}=k]}=1$ if $y_{i}=k$ and 00 otherwise.
$\pi(\theta)\propto\mathcal{L}(\theta)\pi_{0}(\theta).$ (18)
In the following examples, the QOI is interpreted as $\varphi=g$ in the regression problem and as the softmax function of the output, $\varphi=h$, in the classification problem. Rather than considering MSE as in the regression example, the mean centred cross-entropy (KL-divergence) is used as metric between the estimated Bayesian ground truth [^4] Bayesian ground truth for each example is computed by a single SMC-HMC with a large number of samples. $\varphi_{g}$ and the Bayesian estimators $\hat{\varphi}$ in the classification cases as follows
4.3.1 Iris Classification Example
Consider a fully connected deep neural network with only one layer for learning a three-class classification problem on iris data set, that is $D=1$, $n_{0}=4$ and $n_{1}=K=3$, then the dimension of the parameter $\theta$ in neural network is $d=15$. There are $150$ data points in the iris data set, with the first $m=50$ be the training set and the remaining $100$ be the testing set. The precision parameter in likelihood is $\sigma=1$. We take $L=20$, $M=1$, $\Delta t=0.1$ and $M_{0}=I$ in HMC. The convergence results are shown in Figure 10 and crossover point is given in table 3.
4.3.2 MNIST Classification Example
Now, we consider a convolution neural network with two convolution layers followed by two fully connected layers, each convolution layer with $5$ input channels and output channels, $5\times 5$ kernel and stride as $1$. Then the dimensional of parameter $\theta$ is $d=1177$. Using this neural network, we intend to learn a binary classification problem on the MNIST data set by taking the image for "0" and "1". We take the first $m=50$ images for training and following $100$ images for testing. The precision parameter in likelihood is $\sigma=1$. We take $L=20$, $M=1$, $\Delta t=0.001$ and $M_{0}=0.01I$ in HMC. The convergence results are shown in Figure 11 and crossover point is given in table 3.
5 Conclusion
Acknowledgement
Appendix A Techniques for pSMC in practice
The procedure is defined as follows: for core $r=1,...,R$,
$Z^{N,r}$ $=\prod_{j=1}^{J}\sum_{i=1}^{N}\omega_{j}^{i,r}=\prod_{j=1}^{J}\Delta K_{j}^{r}\Delta\hat{Z}_{j}^{r}=K^{r}\hat{Z}^{r},$
where $\Delta K_{j}^{r}=\exp(\max_{i}\log(w_{j}^{i,r}))$ and $K^{r}=\prod_{j=1}^{J}\Delta K_{j}^{r}$; $\Delta{\hat{Z}}_{j}^{r}=\sum_{i=1}^{N}\hat{w}_{j}^{i,r}=\sum_{i=1}^{N}\exp(\log(w_{j}^{i,r})-\max_{i}\log(w_{j}^{i,r}))$ and $\hat{Z}^{r}=\prod_{j=1}^{J}\Delta\hat{Z}_{j}^{r}$. For the small number of importance sampling weights $w_{j}^{i,r}$, we work with log-likelihood for unnormalized weights and modify it with a constant $\hat{w}_{j}^{i,r}$.
Using the same trick again on the normalized constant over $R$ cores, we have, for core $r$,
$Z^{N,r}=K^{r}\hat{Z}^{r}=\exp(\log(\hat{Z}^{r})+\log(K^{r})-\log(K))K,$
where $\log(K)=\max_{r}(\log(\hat{Z}^{r})+\log(K^{r}))$. Since the term $K$ will be cancelled out as we compute the parallel SMC estimator, so we only need the term $\exp(\log(\hat{Z}^{r})+\log(K^{r})-\log(K))$, which could further avoid the computational error due to the small number $K$.
Appendix B Parameterization of $\rho(\theta)$ in QST
  1. 1.
  2. 2.
    Use QR decomposition to get $D=BP$.
  3. 3.
    Create the diagonal matrix $P_{\text{sign}}=\text{diag}(\frac{p_{11}}{|p_{11}|},...,\frac{p_{2^{Q}2^{Q}}}{|p_{2^{Q}2^{Q}}|})$, where $p_{ii}=[P]_{ii}$ for $i=1,2,3,...,2^{Q}$.
  4. 4.
    Construct the unitary matrix $U$ as $U=BP_{\text{sign}}$.
Appendix C Proofs
C.1 Proof relating to Lemma 3.4
Lemma 4.4 Assume Assumption 3.1 and 3.2. Then, for any $J\in\mathbb{N}$, there is a $C>0$ such that for any $N,R\in\mathbb{N}$, suitable $\zeta:\mathsf{\Theta}\rightarrow\mathbb{R}$,
$\mathbb{E}[(F^{N,R}(\zeta)-f(\zeta))^{2}]\leq\frac{C|\zeta|_{\infty}^{2}}{NR}.$
Proof.
$\mathbb{E}[(F^{N,R}(\zeta)-f(\zeta))^{2}]$ $=\mathbb{E}\bigg{[}\bigg{(}\frac{1}{R}\sum_{r=1}^{R}(f^{N,r}(\zeta)-f(\zeta))\bigg{)}^{2}\bigg{]}$
$=\frac{1}{R^{2}}\mathbb{E}\bigg{[}\sum_{r=1}^{R}(f^{N,r}(\zeta)-f(\zeta))^{2}+\sum_{r=1}^{R}\sum_{r^{\prime}=1}^{R}(f^{N,r}(\zeta)-f(\zeta))(f^{N,r^{\prime}}(\zeta)-f(\zeta))\bigg{]}$
$=\frac{1}{R^{2}}\sum_{r=1}^{R}\mathbb{E}[(f^{N,r}(\zeta)-f(\zeta))^{2}]+\frac{1}{R^{2}}\sum_{r=1}^{R}\sum_{r^{\prime}=1}^{R}\mathbb{E}[(f^{N,r}(\zeta)-f(\zeta))]\mathbb{E}[(f^{N,r^{\prime}}(\zeta)-f(\zeta))].$ (19)
By Proposition 3.3, three expectation terms in (19) are expressed as follows
$\mathbb{E}[(f^{N,r}(\zeta)-f(\zeta))^{2}]\leq\frac{C|\zeta|_{\infty}^{2}}{N},\ \mathbb{E}[(f^{N,r}(\zeta)-f(\zeta))]=0.$
Then, we conclude
$\mathbb{E}[(F^{N,R}(\zeta)-f(\zeta))^{2}]\leq\frac{C|\zeta|_{\infty}^{2}}{NR}.$
C.2 Proof relating to Theorem 3.5
Theorem 4.5 Assume Assumption 3.1 and 3.2. Then, for any $J\in\mathbb{N}$, there exists a $C_{\varphi}>0$, which depends on $\varphi$, such that for any $R\in\mathbb{N}$,
[…]
Then the strong scaling in the parallel SMC is described as the MSE defined in (20) decreases as $1/R$ for $R$ nodes, for suitable variables value ($M$,$N$,$J$), with constant time.
Proof.
$\mathbb{E}[(\hat{\varphi}_{\text{pSMC}}-\pi(\varphi))^{2}]$ $=\mathbb{E}\bigg{[}\bigg{(}\frac{F^{N,R}(\varphi)}{F^{N,R}(1)}-\frac{f(\varphi)}{f(1)}\bigg{)}^{2}\bigg{]}$
$=\mathbb{E}\bigg{[}\bigg{(}\frac{F^{N,R}(\varphi)}{F^{N,R}(1)}-\frac{F^{N,R}(\varphi)}{f(1)}+\frac{F^{N,R}(\varphi)}{f(1)}-\frac{f(\varphi)}{f(1)}\bigg{)}^{2}\bigg{]}$
Applying Cauchy-Schwartz inequality, we have
$\leq 2\mathbb{E}\bigg{[}\bigg{(}\frac{F^{N,R}(\varphi)}{F^{N,R}(1)}-\frac{F^{N,R}(\varphi)}{f(1)}\bigg{)}^{2}\bigg{]}+2\mathbb{E}\bigg{[}\bigg{(}\frac{F^{N,R}(\varphi)}{f(1)}-\frac{f(\varphi)}{f(1)}\bigg{)}^{2}\bigg{]}$
$\leq\frac{2|F^{N,R}(\varphi)|_{\infty}^{2}}{|F^{N,R}(1)|^{2}|f(1)|^{2}}\mathbb{E}[(F^{N,R}(1)-f(1))^{2}]+\frac{2}{|f(1)|^{2}}\mathbb{E}\big{[}(F^{N,R}(\varphi)-f(\varphi))^{2}\big{]}$ (21)
Assume Assumption 3.1, there exists a $C_{\varphi}^{\prime}$ such that $\frac{2|F^{N,R}(\varphi)|_{\infty}^{2}}{|F^{N,R}(1)|^{2}|f(1)|^{2}}\leq C_{\varphi}^{\prime}$, and there exists a $C^{\prime}$ such that $\frac{2}{|f(1)|^{2}}\leq C^{\prime}$. Then, following (21), we have
$\mathbb{E}[(\hat{\varphi}_{\text{pSMC}}-\pi(\varphi))^{2}]$ $\leq C_{\varphi}^{\prime}\mathbb{E}[(F^{N,R}(1)-f(1))^{2}]+C^{\prime}\mathbb{E}\big{[}(F^{N,R}(\varphi)-f(\varphi))^{2}\big{]}$
By Lemma 3.4 with $\zeta=\varphi$ and $\zeta=1$ respectively, we have
$\leq\frac{C_{\varphi}^{\prime}C}{NR}+\frac{C^{\prime}C|\varphi|_{\infty}^{2}}{NR}$
$\leq\frac{C_{\varphi}}{NR}.$
Appendix D Figures
D.1 Gaussian cases
@@
Figure 2: Convergence of the Gaussian case for $m=d=8$. In pSMC, the number of mutation steps is $M=10$; in pCN, the number of burn-in samples is $10000$. For both methods, the number of realizations is $100$.
@@
Figure 3: Convergence of the Gaussian case for $m=4$ and $d=16$. In pSMC, the number of mutation steps is $M=10$; in pCN, the number of burn-in samples is $5000$. For both methods, the number of realizations is $100$.
@@
Figure 4: Convergence of the Gaussian case for $m=64$ and $d=16$. In pSMC, the number of mutation steps is $M=d$; in pCN, the number of burn-in samples is $10000$. For both methods, the numbers of realizations is $10$.
@@
Figure 5: Convergence of the Gaussian case for $m=d=32$. In pSMC, the number of mutation steps is $M=d$; in pCN, the number of burn-in samples is $10000$. For both methods, the number of realizations is $10$.
@@
Figure 6: Convergence of the Gaussian case for $m=16$ and $d=64$. In pSMC, the number of mutation steps is $M=d$; in pCN, the number of burn-in samples is $10000$. For both methods, the number of realizations is $10$.
D.2 Quantum State Tomography problem
@@
Figure 7: Convergence of the quantum state tomography example for $m=6$ and $d=16$. In pSMC, the number of mutation steps is $M=d$; in pCN, the number of burn-in samples is 00. For both methods, the number of states is $S=10$.
@@
Figure 8: Convergence of the quantum state tomography example for $m=36$ and $d=64$. In pSMC, the number of mutation steps is $M=d$; in pCN, the number of burn-in samples is 00. For both methods, the number of states is $S=10$.
@@
Figure 9: Convergence of the quantum state tomography example for $m=216$ and $d=256$. In pSMC, the number of mutation steps is $M=d$; in pCN, the number of burn-in samples is 00. For both methods, the number of states is $S=10$.
D.3 Bayesian Neural Network Examples
@@
Figure 10: Convergence of the iris classification example. For both methods, the number of realizations is $10$, $L=20$, $M=1$, $\Delta=0.1$, $M_{0}=I$.
@@
Figure 11: Convergence of the MNIST classification example. For both methods, the number of realizations is $10$, $L=20$, $M=1$, $\Delta t=0.001$, $M_{0}=0.01I$.

### v2

@@ E.2 Details of Gaussian cases
Assume we have $y\in\mathbb{R}^{m}$, $X\in\mathbb{R}^{m\times d}$ and parameter $\theta\in\mathbb{R}^{d}$ connected by the following inverse problem:
@@
Let $X\in\mathbb{R}^{m\times d}$ be a randomly selected full rank matrix and $\sigma=0.01$. The observations are generated as
@@
E.3 Parameterization of $\rho(\theta)$ in QST
  1. 1.
  2. 2.
    Use QR decomposition to get $D=BV$.
  3. 3.
    Create the diagonal matrix $V_{\text{sign}}=\text{diag}(\frac{v_{11}}{|v_{11}|},...,\frac{v_{2^{Q}2^{Q}}}{|v_{2^{Q}2^{Q}}|})$, where $v_{ii}=[V]_{ii}$ for $i=1,2,3,...,2^{Q}$.
  4. 4.
    Construct the unitary matrix $U$ as $U=BV_{\text{sign}}$.
Gaussian m×$\times$×d = 1024
m = 64 > d m = 32 = d m = 16 < d
MSE 8($\pm$ 1)e-07 4($\pm$ 1)e-01 2($\pm$ 0.3)e+01
No.node 1 1 1
Time 2.9e-01 1.0e+00 1.9e+00
E.4 Details of the Bayesian Neural Network
Let weights be $A_{i}\in\mathbb{R}^{n_{i}\times n_{i-1}}$ and biases be $b_{i}\in\mathbb{R}^{n_{i}}$ for $i\in\{1,...,D\}$, we denote $\theta:=((A_{1},b_{1}),...,(A_{D},b_{D}))$. The layer is defined by
@@ E.4 Details of the Bayesian Neural Network
Consider the discrete data set in classification problem, we have $\mathsf{Y}=\{1,...,K\}$ and $n_{D}=K$, then we instead define the so-called softmax function as
@@
Appendix F Figures
F.1 Gaussian cases
@@
Figure 2: Convergence of the Gaussian case for $m=d=8$. In pSMC, the number of mutation steps is $M=10$; in pCN, the number of burn-in samples is $10000$. For both methods, the number of realizations is $R=100$. Error bars is $\pm$±s.e. of the MSE computed in Appendix E.1.
@@
Figure 3: Convergence of the Gaussian case for $m=4$ and $d=16$. In pSMC, the number of mutation steps is $M=10$; in pCN, the number of burn-in samples is $5000$. For both methods, the number of realizations is $R=100$. Error bars is $\pm$±s.e. of the MSE computed in Appendix E.1.
@@
Figure 4: Convergence of the Gaussian case for $m=64$ and $d=16$. In pSMC, the number of mutation steps is $M=d$; in pCN, the number of burn-in samples is $10000$. For both methods, the numbers of realizations is $R=10$. Error bars is $\pm$±s.e. of the MSE computed in Appendix E.1.
@@
Figure 5: Convergence of the Gaussian case for $m=d=32$. In pSMC, the number of mutation steps is $M=d$; in pCN, the number of burn-in samples is $10000$. For both methods, the number of realizations is $R=10$. Error bars is $\pm$±s.e. of the MSE computed in Appendix E.1.
@@
Figure 6: Convergence of the Gaussian case for $m=16$ and $d=64$. In pSMC, the number of mutation steps is $M=d$; in pCN, the number of burn-in samples is $10000$. For both methods, the number of realizations is $R=10$. Error bars is $\pm$±s.e. of the MSE computed in Appendix E.1.
@@
Figure 7: Convergence of the Gaussian case for $m=128$ and $d=32$. In pSMC, the number of mutation steps is $M=d$; in pCN, the number of burn-in samples is $50000$. For both methods, the numbers of realizations is $R=10$. Error bars is $\pm$±s.e. of the MSE computed in Appendix E.1.
@@
Figure 8: Convergence of the Gaussian case for $m=d=64$. In pSMC, the number of mutation steps is $M=d$; in pCN, the number of burn-in samples is $50000$. For both methods, the number of realizations is $R=10$. Error bars is $\pm$±s.e. of the MSE computed in Appendix E.1.
@@
Figure 9: Convergence of the Gaussian case for $m=32$ and $d=128$. In pSMC, the number of mutation steps is $M=d$; in pCN, the number of burn-in samples is $100000$. For both methods, the number of realizations is $R=10$. Error bars is $\pm$±s.e. of the MSE computed in Appendix E.1.
F.2 Quantum State Tomography problem
Figure 10: Convergence of the quantum state tomography example for $m=6$ and $d=16$. In pSMC, the number of mutation steps is $M=d$; in pCN, the number of burn-in samples is 00. For both methods, the number of states is $S=10$. Error bars is $\pm$±s.e. of the MSE computed in Appendix E.1.
Figure 11: Convergence of the quantum state tomography example for $m=36$ and $d=64$. In pSMC, the number of mutation steps is $M=d$; in pCN, the number of burn-in samples is 00. For both methods, the number of states is $S=10$. Error bars is $\pm$±s.e. of the MSE computed in Appendix E.1.
Figure 12: Convergence of the quantum state tomography example for $m=216$ and $d=256$. In pSMC, the number of mutation steps is $M=d$; in pCN, the number of burn-in samples is 00. For both methods, the number of states is $S=10$. Error bars is $\pm$±s.e. of the MSE computed in Appendix E.1.
Figure 13: Convergence of the quantum state tomography example for $m=1296$ and $d=1024$. In pSMC, the number of mutation steps is $M=d$; in pCN, the number of burn-in samples is 00. For both methods, the number of states is $S=10$. Error bars is $\pm$±s.e. of the MSE computed in Appendix E.1.
F.3 Bayesian Neural Network Examples
@@
Figure 14: Convergence of the iris classification example for $m=50$ and $d=15$. For both methods, the number of realizations is $R=10$, $L=20$, $M=1$, $\Delta t=0.1$, $M_{0}=I$. Error bars is $\pm$±s.e. of the loss computed in Appendix E.1.
@@
Figure 15: Convergence of the biMNIST classification example for $m=50$ and $d=1177$ ($n_{1}=n_{2}=5$ and $n_{3}=5$). For both methods, the number of realizations is $R=10$, $L=20$, $M=1$, $\Delta t=0.001$, $M_{0}=0.01I$. Error bars is $\pm$±s.e. of the loss computed in Appendix E.1.
Figure 16: Convergence of the biMNIST classification example for $m=100$ and $d=3587$ ($n_{1}=n_{2}=10$ and $n_{3}=5$). For both methods, the number of realizations is $R=10$, $L=20$, $M=1$, $\Delta t=0.001$, $M_{0}=0.01I$. Error bars is $\pm$±s.e. of the loss computed in Appendix E.1.
Figure 17: Convergence of the biMNIST classification example with $d=3587$ ($n_{1}=n_{2}=10$ and $n_{3}=5$). For both methods, the number of realizations is $R=10$, $L=20$, $M=1$, $\Delta t=0.001$, $M_{0}=0.01I$. The Ground truth here is the data. Error bars is $\pm$±s.e. of the loss computed in Appendix E.1.